OpenStudy (anonymous):

How many solutions does the following system of equations have? x=-1 x^2+y^2=4

jimthompson5910 (jim_thompson5910):

x^2+y^2=4 (-1)^2+y^2=4 ... plug in x = -1 1 + y^2 = 4 y^2 = 4-1 y^2 = 3 I'll let you finish

OpenStudy (anonymous):

Would it be just one solution then?

jimthompson5910 (jim_thompson5910):

no because y^2 = 3 has two solutions (you're forgetting about the plus/minus)

OpenStudy (anonymous):

Oh shoot, thank you!

jimthompson5910 (jim_thompson5910):

its ok, silly mistake

jimthompson5910 (jim_thompson5910):

yw